Minutes of the 5th Committee Meeting of the Dublin University Rifle Club

Moderator Ben Jones

Minutes taker Mark Dennehy

Present Benjamin Jones (Treasurer), Mark Dennehy (Secretary/Match Secretary), Alan Smyth (Chief Range Officer), Conor McDermottroe (Public Relations Officer)

Absent(with apologies) Steven Watterson (Captain), John Keeney (Ammo&Equipment Officer)

Absent Jason Sibley, Deirdre Ninaber (Ladies’ Coordinator)

Date 10 July 2001 1920 – 2150

Distribution DURC Committee

Overview of topics

1 Minutes To-Do List

1 Minutes To-Do List 1 2 Secretarys Report 1 3 Treasurers Report 2 4 CRO Report 2 5 A&E Report 2 6 Offences against the state act (1998) 3 7 Rifle Storage 3 8 Review of the Nationals 4 9 Events Calendar 4 10 DURC 50m Open 4 11 ISC Children’s code of ethics 6 12 Affiliation to the NTSA 6

1 Minutes To-Do List

The todo list from the previous meeting was examined and status of the various jobs reported on, most being completed. Mark still has to send off letters regarding the DURC Open.

2 Secretarys Report

Mark reported on the correspondance received by the club. The Trinity Top Ten now has 33 entries from outside the club, and so will be self-financing. DURC entries have still not been arranged. Ben said he would enter the Air Rifle league; Conor McDermottroe, Peter Byrne and Alan Smyth will enter the 25yd league and Peter and Alan will enter the 50m league. Mark is to continue to chase down other DURC entries and send off the stickers to outside entrants. Task: Arrange DURC entries into the TTT (Mark)

Sponsorship requests were discussed, Eircom having turned us down. Mark said we could resubmit in October once the Eircom sale has finalised. Conor recommended not requesting a specific sum as Eircom would not give smaller amounts if they had them if

3 Treasurers Report

we gave a specific sum they didn’t have. Asking for prizes was also mentioned as a good idea.

A discussion of the Competition Attendance Report distribution list sprang up, Alan concerned it was becoming a “spam list”. Mark pointed out that this was the idea – to raise the level of awareness of the club in the people on the list. Mark also stated that Cathy has reported that she thinks the CARs are a great idea as she is now kept up-to-date with club happenings. It was agreed to maintain the CAR distribution.

Entry fees for Air nationals were also discussed, with debate over responsibility for Mark being incorrectly entered. It was felt by most that the NTSA were to blame. However, Mark felt the onus was on him to check: it was concluded not to bug the NGB about this and deduct the amount from Mark’s copious credits.

3 Treasurers Report

Ben reported that Cathy may give the money originally slated for hiring the WTSC range for the colours to the Bisley trip instead, especially as he had made the Centenary Postal finals. Given the sale of the lead on the range (Mark congradulated Alan on a job well done) and money received for the Nationals, the bank account now stands at £230 but the bill for the Nationals is reportedly en route along with one from RRPC from a previous competition.

The matter of the payment of the Indoor Nationals bill arose - Mark said he was going to point out in the strongest terms to the NTSA that we had paid the bill and they had the money in their account when the comment was entered in the NTSA’s minutes. Task: Complain to the NTSA about defaming our name regarding bills. (Mark)

4 CRO Report

Alan reported the new halogen floodlamp was installed and he would appreciate comment on the positioning of same. Applications for removal orders had been made and received for the Squad training and the DURC Open. The lead on the range had been sold for £72.

5 A&E Report

John was not present and so no report was given.

6 Offences against the state act (1998)

Conor has conferred with officials from the Dept. of Justice and we are exempt by previous OATS Acts, the exemption is thus contained in section 12 sub-section 4 of the 1998 amendment. Thus we do not violate the Offences against the State act five nights a week.

7 Rifle Storage

7 Rifle Storage

Mark gave a brief account of the events following the meeting between John Keeney, Pat Moray and Tim Cooper. Originally to arrange the ammunition accounting procedures, Pat then sprang the demand to remove all privately held rifles from the College and refused to negotiate or further discuss the matter, either at the meeting or later in private with John. Mark reported that a meeting with John, Pat and Terry McAuley, Director of Sport, was arranged for next week. Mark also reported that in light of the circumstances he was keeping his new rifles at home for the moment.

Mark pointed out that while Pat had no legal ability to seize rifles, no liability for privatly owned rifles and no work imposed on him by their presence, we had a serious problem on our hands and treading cautiously was necessary. John was the point man for this set of negotiations and the rest of us should steer clear of Pat for the moment.

Options were discussed. Buying a new gunsafe for privately owned rifles only was suggested; Mark reported John thought that that would not work with Pat. It was pointed out at the same time that a new gunsafe was needed on the range anyway to take the remaining rifles from Front Arch to the range. The accounting problems with the ammunition were seen as the most likely source of Pat’s irritation. It was suggested by Mark that it was likely that Pat had picked what Pat had thought was a minor irritation to us as a punitive measure, but he had in fact chosen a measure that was far too useful for us to give up.

The final consensus was that a serious honest discussion had the ability to sort the mess out as fast as possible. If not, we had a very good case to take to the Provost, in that we have been in college since 1840 (since 1962 in the modern form); Pat has never had objections prior to this; The Gardai have assured us they have no problems with the arrangements whatsoever; Every shooter in DURC with a privately owned rifle is either a National Champion or (in Steven’s case) a very experienced shooter who would be a National Champion if he wasn’t from another country. Mark also pointed out that the facility to allow experienced members to own their own rifles was too important to give up as it allows new shooters to invest in their own equipment with a minimum of problems. It was also pointed out that backing off the issue would lead to further measures, such as Pat pointing out that Air rifles were safer as they can’t shoot through the ceiling of the range into the paint shop overhead and that therefore he would no longer allow .22s in college.

8 Review of the Nationals

The “buddy” system proved to be excellent, with Mark, Conor and Alan testifying to it’s usefulness. (John had earlier agreed with this appraisal). It was agreed to keep it as the standard procedure for competitions.

The scopes were mentioned as being unsuitable for 50m, several are broken and do not focus well; all are useful at 25yds but near-useless at 50m with shot holes being almost invisible. Alan agreed to try to fix some of the scopes. Mark reported an interesting

9 Events Calendar

phenonomen with his scope, in that it imparted a 45 degree rotation to the image through an unknown mechanism. Task: Attempt repair of club scopes. (Alan)

The Code of Conduct document had not been prepared in time for the Nationals due to Steven’s departure for the Island Games. Mark agreed to prepare it for the DURC Open. Task: Prepare code of conduct for DURC before Open. (Mark)

The pre-competition team talk was mentioned and accepted as a good idea. It was agreed that it should be held as standard procedure for people attending competitions with DURC, usually on the friday beforehand on the range. Topics to mention in the talk were listed as: the course of fire for the competition; characteristics of the range; logistical arrangements; signing of the code of conduct and discussion of same; techniques for shooting in wind if appropriate; and any other pertinant topics. Alan was concerned that the level of information might be too shallow, a little knowlege being a dangerous thing. Mark retorted that no information was a lot more dangerous. The level of information was agreed on as being something that required observation.

9 Events Calendar

Deferred due to the lack of a college almanack for the next academic year.

10 DURC 50m Open

Mark reported that Shilen have donated a voucher for a turned barrel blank for a prize. It was agreed this was an excellent prize and it was put up as first prize in the championship. The entry fees were discussed and set at £15 per person for entering both the english match and the rapid-fire and the barbecue. It was estimated that 20 such entrants would be the break-even point for the event; DURC entrants not included as DURC members pay £5 apiece. It was agreed to charge the sporter people £3 each for the barbecue on an informal basis on the day. Thus the total budget for the event is £300. Task: Confirm NTSA levy amounts. (Mark)

Publicising the competition was discussed, Mark reported that it was in the NTSA events calendar, on the NTSA website, that it had been announced at the Nationals during the Prone and 3P matches, and that it was also in the current Marksman which had gone to press that morning. Sending a personal invitation to shooters was considered and deemed a good idea, as 30 letters would cost £10 and one entrant would pay for it and more. Mark suggested writing to N.Ireland shooters, mentioning the first prize and an offer to loan club rifles/barrels for the competition to these shooters to get around the problems with licencing. This was deemed an excellent idea. Task: Write to (and phone) clubs to announce the DURC Open (Mark) Task: Send personal notices of the DURC Open to various shooters (Mark)

The option of McRae handicap scores for the rapid-fire competition was mentioned by Mark but felt to be a bad idea as the handicaps were based on slow-firing not rapid-firing

10 DURC 50m Open

and so the option was dismissed.

The logistics of the barbecue were discussed. Mark announced he had notes on vegetarian options, and he went through the notes from last year’s BBQ. He also pointed out that we had gone over a budget last year before he was informed of the budget and that he was tired of being reminded at every turn as it was unfair for one person to take responsibility for doing a task and everyone else to then criticise them when they did no work at all. The menu for this year’s BBQ was discussed and a budget of £100 agreed on for the food. Conor proved to have some culinary talent and was recruited as deputy chef, Mark pointing out that he had gotten sick from sunstroke last year as a result of not having someone to take turns at the BBQ last year. Task: Arrange BBQ. (Mark)

The prizes were then extensivly discussed and it was decided that: First place in the championship would get the shilen barrel and the shooter’s name would be engraved on a perpetual shield, to be displayed in the shooter’s club for the year. Second and third places were to be picked on a shopping trip to the trophy shop in Frederick St. The rapid-fire competition would use wine as prizes. Conor volunteered to arrange this due to his contacts in O’Briens. Sponsorship was mentioned at this prospect. Wine was also to be the prizes for class A. The class B prizes were to be the remaining crystal from the DURC air open; class C was to be marble or other plaques, and class D was to be either medals or plaques depending on stock available in Frederick St.

Mark pointed out at this stage that we should invest in our own trophy centers next year. Alan also announced that if he managed to get funding for a rapid prototyping machine for his postgraduate work, that he would make next year’s prizes.

Mark suggested a teams approach as in the Nationals - that the best 3 in each club have their scores tracked and the best club announced but not with prizes. This was agreed, with Conor suggesting we massage the figures to put DURC in the best light, by announcing averages or total gunscore or whatever.

Mark announced he would talk to Nick West about the classification scheme, which has problems in that it is classifying outdoor shoots based on indoor results, which is unfair. Task: Appeal classification scheme to NTSA. (Mark)

Ben and another volunteer were to go on a shopping trip for prizes in Frederick St with a budget of £100. Task: Go shopping for prizes in Frederick St. (Ben&Volunteer)

Logistics were then discussed. It was agreed that Mark, being closest to FSC, would drive out to FSC on sunday morning with the BBQ, food, his laptop computer, the targets, cashbox, guest book and his rifle&gear, so that the competition could begin by 0930.

Alan and Ray (and maybe Peter Byrne) would drive out from DURC later with the rest of the rifles and equipment, along with the prizes, guages and ammo boxes for the rapid-fire in case any competitors had none - it being agreed to use the 10-round ammo boxes for safety reasons. Ben agreed to help score targets, but it was agreed that a second volunteer was needed.

1600 in DURC on the friday before the competition was set as the time for people

11 ISC Children’s code of ethics

to pre-pack and verify they had all of their gear. People packing their own bags (Ray, Alan, Mark, Steven) were to leave lists of the gear they had packed so that other people could cross-check instead of searching the range for gear that was allready packed.

11 ISC Children’s code of ethics

Mark reported that the ISC had released their code of ethics for children in sport and that it applied to us and required us to make constitutional changes and appoint two new officers - a children’s officer and a liason officer between the club and the health board and other statuory bodies. It was agreed to appoint these when it became necessary. Task: Find out more details regarding ISC’s new code of ethics. (Mark)

12 Affiliation to the NTSA

Mark also reported that discussion had taken place in the last NTSA meeting regarding the new policy of individual membership of the NTSA and their umbrella insurance policy. A discussion about insuring items twice took place, Mark pointing out that it was legal if the two policies applied in different physical areas (ie, inside and outside college). Mark stated he felt the individual membership idea would go ahead as we didn’t have to clout to stop it; that it would probably be £5 per student and that it would be possible to arrange that the first competition did not require membership. This was accepted; though Alan wondered if we needed to pay this fee for our members. Mark said UCD did not, but he felt we should as we can’t transfer the money from one grant area to another easily while UCD can. Ben asked about the billing arrangements - the start or end of year or in dribs and drabs? And could a flat fee be negotiated? Mark agreed to find out. The inevitable query of “what does the NTSA do for us” was asked and Mark tried to outline points - the basic one being that without the NTSA we don’t have a national sport or a national squad. Task: Negotiate a good arrangement on behalf of DURC in affiliations. (Mark)

T.B.N.O.B.T.M.W.D.C.

Steven Watterson, Club Captain

Mark Dennehy, Club Secretary

12 Affiliation to the NTSA

Tasks agreed on at meeting

Mark Arrange DURC entries into the TTT (10 July 2001) [/141] . . . . . /2 Mark Complain to the NTSA about defaming our name regarding bills.

(10 July 2001) [/183] . . . . . . . . . . . . . . . . . . . . . . . . . . . . /2

Alan Attempt repair of club scopes. (10 July 2001) [/267] . . . . . . . . . /2 Mark Prepare code of conduct for DURC before Open. (10 July 2001) [/273] /2 Mark Confirm NTSA levy amounts. (10 July 2001) [/306] . . . . . . . . . /2 Mark Write to (and phone) clubs to announce the DURC Open (10 July

2001) [/322] . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . /2

Mark Send personal notices of the DURC Open to various shooters (10

July 2001) [/324] . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . /2

Mark Arrange BBQ. (10 July 2001) [/345] . . . . . . . . . . . . . . . . . . /2 Mark Appeal classification scheme to NTSA. (10 July 2001) [/379] …. /2 Ben&VolunteerGo shopping for prizes in Frederick St. (10 July 2001) [/384] …. /2 Mark Find out more details regarding ISC’s new code of ethics. (10 July

2001) [/417] . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . /2

Mark Negotiate a good arrangement on behalf of DURC in affiliations. (10 July 2001) [/442] . . . . . . . . . . . . . . . . . . . . . . . . . . . . /2